OpenStudy (zappy620):

Find the derivative of cosx/(1+sinx)

OpenStudy (zappy620):

The answer is \[\frac{ 1 }{1 -sinx }\]

OpenStudy (zappy620):

The answer I got was \[\frac{ -1 }{ 1+sinx}\]

OpenStudy (zappy620):

My question is are those the same thing?

OpenStudy (holsteremission):

They are not. \[\frac{1}{1-\sin x}=\frac{-1}{-(1-\sin x)}=\frac{-1}{-1+\sin x}\neq\frac{-1}{1+\sin x}\]

OpenStudy (zappy620):

Oh ok

OpenStudy (zappy620):

So its a mistake to factor out a -1 in the denominator in this situation? \[\frac{ -sinx -1}{ (sinx+1)^2 }\]

OpenStudy (zappy620):

i mean numerator*

OpenStudy (holsteremission):

Your answer is correct.\[\begin{align*} \frac{\mathrm d}{\mathrm dx}\left[\frac{\cos x}{1+\sin x}\right]&=\frac{\dfrac{\mathrm d}{\mathrm dx}[\cos x](1+\sin x)-\dfrac{\mathrm d}{\mathrm dx}[1+\sin x]\cos x}{(1+\sin x)^2}\\[1ex] &=\frac{-\sin x(1+\sin x)-\cos^2x}{(1+\sin x)^2}\\[1ex] &=\frac{-\sin x-(\sin^2x+\cos^2x)}{(1+\sin x)^2}\\[1ex] &=-\frac{\sin x+1}{(1+\sin x)^2}\\[1ex] &=-\frac{1}{1+\sin x} \end{align*}\]

OpenStudy (zappy620):

Should I type out the work I have if it's confusing to understand what i'm asking?

OpenStudy (zappy620):

Oh really? The derivative calculator had a different answer and I was confused as to why.

OpenStudy (holsteremission):

Yup. You can always integrate to check. \[\begin{align*} -\int\frac{\mathrm dx}{1+\sin x}&=-\int\frac{1-\sin x}{\cos^2x}\,\mathrm dx\\[1ex] &=\int\left(\sec x\tan x-\sec^2x\right)\,\mathrm dx\\[1ex] &=\sec x-\tan x+C\\[1ex] &=\frac{1-\sin x}{\cos x}+C\\[1ex] &=\frac{1-\sin^2x}{\cos x(1+\sin x)}+C\\[1ex] &=\frac{\cos^2x}{\cos x(1+\sin x)}+C\\[1ex] &=\frac{\cos x}{1+\sin x}+C \end{align*}\]

OpenStudy (mathstudent55):

\(\dfrac{d}{dx}\dfrac{\cos x}{1 + \sin x} =\) \(=\dfrac{d}{dx}[\cos x(1 + \sin x)^{-1}]\) \(=\cos x \dfrac{d}{dx}(1 + \sin x)^{-1} + (1 + \sin x)^{-1} \dfrac{d}{dx} \cos x\) \(= \cos x [-(1 + \sin x)^{-2}]\cos x + (1 + \sin x)^{-1}(- \sin x)\) \(= -\dfrac{\cos ^2 x}{(1 + \sin x)^2} - \dfrac{\sin x}{1 + \sin x} \) \(= -\dfrac{\cos ^2 x}{(1 + \sin x)^2} - \dfrac{\sin x(1 + \sin x)}{(1 + \sin x)^2} \) \(= \dfrac{-\cos ^2 x - \sin x - \sin^2 x}{(1 + \sin x)^2}\) \(= -\dfrac{\sin^2 x + \cos ^2 x + \sin x}{(1 + \sin x)^2}\) \(= -\dfrac{1 + \sin x}{(1 + \sin x)^2}\) \(= -\dfrac{1 }{1 + \sin x}\)

OpenStudy (zappy620):

Alright thanks guys :D

OpenStudy (mathstudent55):

You're welcome.